# spacing=nonfrench

### v2

@@ Abstract
Emerging AI accelerators increasingly adopt wafer-scale manufacturing technologies, integrating hundreds of thousands of AI cores in a mesh-based architecture with large distributed on-chip memory (tens of GB in total) and ultra-high on-chip memory bandwidth (tens of PB/s). However, current LLM inference systems, optimized for shared memory architectures like GPUs, fail to fully exploit these accelerators.
@@
Evaluations show that WaferLLM achieves 200$\times$ better wafer-scale accelerator utilization than state-of-the-art systems. On a commodity wafer-scale accelerator, WaferLLM delivers 606$\times$ faster and 22$\times$ more energy-efficient GEMV compared to an advanced GPU. For LLMs, based on 16-bit data type, WaferLLM achieves 2700 toks/sec/req decode speed on Llama3-8B model and 840 toks/sec/req decode speed on Qwen2-72B model, which enables 39$\times$ faster decoding with 1.7$\times$ better energy efficiency. We anticipate these numbers will grow significantly as wafer-scale AI models, software, and hardware continue to mature.
@@ 1 Introduction
Large Language Model (LLM) inference is a rapidly growing workload. It has two phases [14]: (i) the prefill phase, which processes input tokens (the prompt) and spends most of its cycles on General Matrix Multiply (GEMM); and (ii) the decode phase, which generates tokens one by one in an autoregressive manner, primarily performing General Matrix-Vector Product (GEMV). Decode requires repeatedly loading the entire LLM model into on-chip memory, with GEMV dominating its cycles. Since LLMs generate many tokens, inference is constrained by GEMV latency, making it inherently memory-bandwidth-bound.
To address memory bandwidth bottlenecks, AI accelerators are increasingly adopting system-on-wafer integration [20]. This approach scales chip area to a full wafer, up to 100$\times$ larger than a typical GPU die, enabling significantly more on-chip cores, memory and bandwidth. Examples include Cerebras WSE [23] and upcoming Tesla Dojo [38]. The Cerebras WSE-2, for instance, integrates 850,000 cores with 40GB of on-chip memory—1000$\times$ more than GPUs—and provides 22PB/s memory bandwidth, 7000$\times$ higher than GPUs. TSMC predicts widespread adoption of system-on-wafer integration due to its performance advantages, energy efficiency in connecting dies, and lowering cost, with IEEE forecasting a wave of wafer-scale computers by 2027 [20].
@@
The above idea drives the design of WaferLLM, the first wafer-scale LLM inference system, yielding several contributions:
@@
(1) Device model for wafer-scale accelerators. We propose the PLMR model [^1] PLMR model can be pronounced as “Plummer” , which captures the following hardware properties of wafer-scale accelerators: (i) Massive Parallel cores (P): Millions of cores can be integrated on a large wafer, requiring systems to effectively partition LLMs and their operations. (ii) Highly non-uniform memory access Latency (L): Inter-core data access exhibits significant variation, with latency differences up to 1000$\times$, necessitating system to mitigate this. (iii) Constrained local Memory (M): Each core has limited memory (tens of KBs to several MBs), requiring efficient memory usage. (iv) Limited hardware-assisted Routing (R): The NoC routing hardware supports small messages (e.g., a few bytes), with headers and address encoding limited to a few bits, restricting routing paths. Careful communication path planning is crucial to avoid falling back to slower software-based routing.
@@
(2) Wafer-scale LLM parallelism. We propose an effective LLM parallelism method for wafer-scale accelerators, fully compliant with the PLMR model. In the prefill phase, we design fine-grained partitioning to achieve million-core parallelism. For the decode phase, where tensor dimensions are insufficient for partitioning, we design fine-grained replication to enable parallelism with minimal communication costs. As a result, WaferLLM achieves larger-scale and finer-grained parallelism (satisfying P in PLMR) than GPU-based approaches. Additionally, we replace conventional GPU-based GEMM and GEMV operators in existing LLM models with new versions designed for the PLMR model (satisfying L, M and R) and propose tensor placement strategies that eliminate matrix transpositions, costly with a mesh NoC (satisfying L).
@@
We also design a scalable KV-cache management method for wafer-scale devices. This approach features a novel KV cache shift method to ensure balanced core usage (satisfying P and M), avoiding skewed utilization of cores caused by KV cache concatenation methods common on GPUs.
@@
(3) Wafer-scale GEMM. We propose MeshGEMM, a scalable GEMM algorithm for wafer-scale accelerators, enabling WaferLLM to fully accelerate its prefill phase. Unlike conventional distributed GEMM algorithms, MeshGEMM achieves full PLMR compliance by leveraging two key operations: cyclic shifting and interleaving. Cyclic shifting ensures algorithm correctness while maintaining bounded usage of local memory (satisfying M) and routing resources (satisfying R). The interleaving operation minimizes multi-hop communication in the mesh NoC, effectively reducing the overhead of highly non-uniform memory latency (satisfying L).
@@
(4) Wafer-scale GEMV. We propose MeshGEMV, a scalable GEMV algorithm for wafer-scale devices, enabling WaferLLM to effectively accelerate its decode phase. Unlike existing GEMV implementations, MeshGEMV uses a novel two-way K-tree allreduce algorithm to aggregate local GEMV results across massive cores. This algorithm ensures minimal routing resource usage (satisfying R) and reduced communication paths (satisfying L).
@@
We implemented WaferLLM on the Cerebras WSE engine using approximately 7,000 lines of CSL (a C-like programming language) for LLM parallelism, MeshGEMM, and MeshGEMV, and 2,000 lines of Python for loading LLM checkpoints and launching inference.
@@
We conducted end-to-end LLM inference experiments with various models, including full LLaMA3-8B and LLaMA3-13B, as well as subsets of layers of LLaMA3-70B, CodeLLaMA-34B, and QWen2-72B. By combining wafer-scale LLM parallelism, GEMM and GEMV, WaferLLM outperforms state-of-the-art (SOTA) systems: (i) 100-200$\times$ faster than T10 [24], the SOTA system for massive cores with a distributed on-chip memory architecture, and (ii) 200-400$\times$ faster than Ladder [41], the SOTA system for shared-memory architectures.
@@
Micro-benchmarks further show that MeshGEMM is 2–3$\times$ faster than SUMMA [39], the default optimized GEMM for Cerebras WSE, and Cannon [7], the SOTA GEMM for supercomputers with large-scale mesh architectures. MeshGEMV achieves 4–8$\times$ speedups over Cerebras’s optimized GEMV. Additionally, WaferLLM ’s cache shift method is up to 400$\times$ more scalable than the KV cache SOTA on GPUs, such as PagedAttention [19].
@@
Combining these benefits, WaferLLM (using Cerebras WSE-2) outperforms vLLM (using A100) by 606$\times$ in GEMV operations and achieves 22$\times$ better energy efficiency. This comparison is fair, as both WSE-2 and A100 are manufactured using TSMC’s 7nm process. For full LLM inference, WaferLLM delivers a 38$\times$ faster decode rate (tokens/s) and is 1.7$\times$ more energy-efficient (token/J) than vLLM. The reduced gains from GEMV to LLM are due to current limitations in Cerebras’s software, hardware, and existing LLM model designs. We anticipate stronger performance as wafer-scale AI computing matures and these limitations are addressed.
@@ 2.1 LLM inference and its key constraint
An LLM inference system typically performs auto-regressive token-by-token generation, as illustrated in Figure 1. The model comprises multiple transformer layers, dominated by self-attention and feedforward blocks. Inference operates in two phases: prefill and decode. The total cycles of the prefill phase are dominated by GEMM operations (shown by 1 ). Similarly, the total cycles of the decode phase are dominated by GEMV operations (shown by 2 ).
LLM inference is memory bandwidth-bound. Model weights (10–100 GB) must be repeatedly fetched from external memory during inference, as GPUs typically have only  100 MB of on-chip memory. Generating a single token requires transferring tens of GBs, and producing thousands of tokens per second demands hundreds of TB/s bandwidth, far exceeding the capabilities of HBM on current GPUs.
While tensor parallelism across GPUs can increase bandwidth, mitigating communication bottlenecks in a large GPU cluster remains challenging. Also, adding GPUs improves throughput for concurrent queries but does not reduce response time, as each query is still memory bandwidth-limited.
@@ 2.2 Reasons for wafer-scale accelerators
Chip Area 400 $\sim$ 800 mm2 30,000 $\sim$ 160,000 mm2 [38]
# Transistors 10s Billions Trillions
# Cores 1,000 $\sim$ 10,000s 100,000 $\sim$ 1,000,000s
On-Chip Memory 100s MBs 10s GBs
Memory Bandwidth TBs/s 10s PBs/s
Attached HBM 10s GBs TBs (via TSMC SoW)
Die-to-Die Bandwidth 100s GBs/s (via PCB) 10s TB/s (via Wafer)
Die-to-Die Power 10s pJ/bit (via PCB) 0.1s pJ/bit (via Wafer)
@@
Performance advantages. System-on-wafer technology allows trillions of transistors to be integrated into a single wafer-scale chip—100$\times$ more than a typical GPU die, shown in Table 1. This enables millions of AI-optimized cores, providing tens of GBs of on-chip memory and up to tens of PB/s memory bandwidth—1,000$\times$ higher than a standard GPU’s several TB/s. Future wafer-scale chips can also attach 40–80$\times$ more HBM chips to their edge compared to a standard die [20].
Integration efficiency. System-on-wafer excels at integrating massive parallel cores, with wafer-based die-to-die connections offering up to 10$\times$ more bandwidth per unit area and nearly 100$\times$ better power efficiency per bit than conventional PCB-based I/O (e.g., NVIDIA NVLink), as shown in Table 1.
Lower cost. Wafer-scale integration can lower the manufacturing cost, since the significant fraction of the cost of fabrication (typically 30-50%) is related to testing and packaging the individual chips [44]. Additionally, wafer-scale integration has made notable progress in yield improvement. Companies such as TSMC are also developing techniques to integrate fully tested dies on a single wafer, further enhancing yield.
@@ 2.3 Challenges for wafer-scale LLM inference
The massive-scale mesh architecture presents challenges for several LLM operations due to their high data movement demands: (i) managing LLM models and KV cache, (ii) GEMM operations during the prefill phase, and (iii) GEMV operations during decoding. Other operations, such as element-wise computations such as dot-product and activation functions, require no data movement and naturally benefit from parallelism. Operations needing allreduce, such as RMSNorm and Softmax, can leverage GEMV solutions.
@@ 3.1 The PLMR model
    Massive Parallelism (P): A wafer-scale accelerator can easily be equipped with millions of parallel cores, compared to thousands in GPUs. Each core features a local hardware pipeline that overlaps data ingress, egress, computation, and memory access at the cycle level. This requires the computation to be partitioned at a massive scale and a fine-grained schedule to overlap computation, memory access, and NoC communication.
@@
    Highly non-uniform memory access Latency (L): Accessing memory on other cores in a mesh exhibits highly non-uniform latency. In a mesh with $N_{w}\times N_{h}$ cores, the maximum NoC hops to a remote core is $\max(N_{w},N_{h})$. For a million-core mesh, this can reach 1000 hops, causing a 1000$\times$ latency difference between local and remote memory access. Therefore, it is crucial for the computation to minimize long-range communication whenever possible.
@@
    Constrained local Memory (M): Each core has a small local memory (tens of KBs to several MBs), as performance and energy efficiency decline with larger capacities [43]. As a result, computation data must be explicitly partitioned into fine-grained chunks to fully fit within the constraints of each core’s local memory.
@@
    Constrained Routing resources (R): The message size in the NoC of a wafer-scale accelerator is extremely limited (e.g., a few bytes). This constraint requires message headers (e.g., address encoding) to be restricted to just a few bits, maximizing the capacity for actual data transfer. Consequently, only limited routing paths can be used, and the software system must carefully plan these paths.

### v3

# WaferLLM: Large Language Model Inference at Wafer Scale

@@ Abstract
Emerging AI accelerators increasingly adopt wafer-scale manufacturing technologies, integrating hundreds of thousands of AI cores in a mesh architecture with large distributed on-chip memory (tens of GB in total) and ultra-high on-chip memory bandwidth (tens of PB/s). However, current LLM inference systems, optimized for shared memory architectures like GPUs, fail to exploit these accelerators fully.
@@
Evaluations show that WaferLLM achieves up to 200$\times$ higher accelerator utilization than state-of-the-art methods. Leveraging a wafer-scale accelerator (Cerebras WSE2), WaferLLM delivers GEMV operations 606$\times$ faster and 16$\times$ more energy-efficient than on an NVIDIA A100 GPU. For full LLM inference, WaferLLM achieves 10–20$\times$ speedups over A100 GPU clusters running SGLang and vLLM. These advantages are expected to grow as wafer-scale AI models, software, and hardware continue to mature. WaferLLM is open-sourced at https://github.com/MeshInfra/WaferLLM.
@@ 1 Introduction
Large Language Model (LLM) inference is a rapidly growing workload. It has two phases [15]: (i) the prefill phase, which processes input tokens (the prompt) and spends most of its cycles on General Matrix Multiply (GEMM); and (ii) the decode phase, which generates tokens one by one in an autoregressive manner, primarily performing General Matrix-Vector Product (GEMV). Decode requires repeatedly loading the entire LLM model into on-chip memory, with GEMV dominating its cycles. Since LLMs generate many tokens, especially in the test-time scaling scenario, such as the OpenAI-o1/o3 [16, 33] and DeepSeek-R1 [13], inference is constrained by GEMV latency, making it inherently memory-bandwidth-bound.
To address memory bandwidth bottlenecks, AI accelerators are increasingly adopting system-on-wafer integration [21]. This approach scales chip area to a full wafer, up to 100$\times$ larger than a typical GPU die, enabling significantly more on-chip cores, memory, and bandwidth. Examples include Cerebras WSE [24] and upcoming Tesla Dojo [41]. The Cerebras WSE-2, for instance, integrates 850,000 cores with 40GB of on-chip memory, 1,000$\times$ more than GPUs, and provides 22PB/s memory bandwidth, 7,000$\times$ higher than GPUs. TSMC anticipates widespread adoption of system-on-wafer integration, citing advantages in performance, energy-efficient die-to-die communication, and cost reduction. IEEE similarly forecasts a surge in wafer-scale computing by 2027 [21]. Wafer-scale accelerators are already seeing real-world deployment, particularly in model serving. In February 2025, Mixtral and Perplexity adopted wafer-scale chips, achieving cost parity with GPUs in terms of tokens per dollar [40, 44]. G42 now operates data centers fully outfitted with wafer-scale accelerators, and Cerebras has secured major commercial contracts [39].
@@
The above idea motivates WaferLLM, the first wafer-scale LLM inference system, yielding several contributions:
@@
(1) Device model for wafer-scale accelerators. We propose the PLMR model [^1] PLMR model can be pronounced as “Plummer” , which captures the key hardware properties of wafer-scale accelerators: (i) Massive Parallel cores (P): Millions of cores can be integrated on a large wafer, requiring systems to effectively partition LLMs and their operations. (ii) Highly non-uniform memory access Latency (L): Inter-core data access exhibits significant variation, with latency differences up to 1,000$\times$, necessitating the system to mitigate this. (iii) Constrained per-core local Memory (M): Each core has limited memory (tens of KBs to several MBs), requiring efficient memory usage. (iv) Limited hardware-assisted Routing (R): The NoC routing hardware is constrained by the area size per core and can only support a limited number of routing paths, e.g., less than $2^{5}$ on Cerebras WSE-2.
@@
(2) Wafer-scale LLM parallelism. We propose an effective, PLMR-compliant LLM parallelism policy for wafer-scale accelerators. In the prefill phase, we design fine-grained partitioning to achieve million-core parallelism. For the decode phase, where tensor dimensions are insufficient for partitioning, we design fine-grained replication to enable parallelism with minimal communication costs. As a result, WaferLLM achieves larger-scale and finer-grained parallelism (satisfying P in PLMR) than GPU-based approaches. Additionally, we replace conventional GPU-based GEMM and GEMV operators with new algorithm designed for the PLMR model (satisfying L, M, and R) and propose tensor placement strategies that eliminate matrix transpositions, which are costly with a mesh NoC (satisfying L).
@@
We also designed a scalable KV-cache management method for wafer-scale devices. This approach features a novel KV cache shift method to ensure balanced core usage (satisfying P and M), avoiding skewed utilization of cores caused by KV cache concatenation methods common on GPUs.
@@
(3) Wafer-scale GEMM. We propose MeshGEMM, a scalable GEMM algorithm for wafer-scale devices, accelerating the prefill phase. Unlike conventional distributed GEMM algorithms, MeshGEMM achieves full PLMR compliance by leveraging two key operations: cyclic shifting and interleaving. Cyclic shifting ensures algorithm correctness while maintaining bounded usage of local memory (satisfying M). The interleaving operation minimizes communication latency in the mesh NoC, effectively reducing the overhead of highly non-uniform memory latency (satisfying L) and routing resources (satisfying R).
@@
(4) Wafer-scale GEMV. We propose MeshGEMV, a scalable GEMV algorithm for wafer-scale devices, accelerating the decode phase. Unlike existing GEMV implementations, MeshGEMV uses a novel K-tree allreduce algorithm to aggregate local GEMV results across massive cores. This algorithm ensures routing resource usage meets the hardware’s limitation (satisfying R) and reduces communication latency (satisfying L).
@@
We implemented WaferLLM on the Cerebras WSE engine using approximately 7,000 lines of CSL (a C-like programming language) for LLM parallelism, MeshGEMM, and MeshGEMV, and 2,000 lines of Python for loading LLM checkpoints, launching inference, and execution parallelism policy.
@@
We conducted end-to-end LLM inference experiments with various models, including full LLaMA3-8B and LLaMA2-13B, as well as subsets of layers of CodeLLaMA-34B, and QWen2-72B. By combining wafer-scale LLM parallelism, GEMM and GEMV, WaferLLM outperforms state-of-the-art (SOTA) systems: (i) 100-200$\times$ faster than T10 [25], the SOTA system for massive cores with a distributed on-chip memory architecture, and (ii) 200-400$\times$ faster than Ladder [45], the SOTA system for shared-memory architectures.
@@
Micro-benchmarks further show that MeshGEMM is 2-3$\times$ faster than SUMMA [42], the default optimized GEMM for Cerebras WSE, and Cannon [6], the SOTA GEMM for supercomputers with large-scale mesh architectures. MeshGEMV achieves 4-8$\times$ speedups over Cerebras’s optimized GEMV, and 606$\times$ than a single A100 GPU. Additionally, WaferLLM ’s cache shift method is up to 400$\times$ more scalable than the KV cache SOTA on GPUs, such as PagedAttention [20].
@@
Combined, WaferLLM (on Cerebras WSE-2) outperforms SGLang (on single A100) by 30-40$\times$. Compared to the optimal performance of SGLang on A100 multi-GPUs connected with NVLink and RDMA, WaferLLM delivers a 10-20$\times$ faster end-to-end speed and is 2.5$\times$ more energy-efficient. The reduced gains from GEMV to LLM are due to current limitations in software, hardware, and existing LLM model designs. We anticipate stronger performance as wafer-scale AI computing matures and these limitations are addressed.
@@ 2.1 LLM inference and its key constraint
An LLM inference system typically performs auto-regressive token-by-token generation, as illustrated in Figure 1. The model comprises multiple transformer layers, dominated by self-attention and feedforward blocks. Inference operates in two phases: prefill and decode. The total cycles of the prefill phase are dominated by GEMM operations (shown by 1 ). While the total cycles of the decode phase are dominated by GEMV operations (shown by 2 ).
LLM inference is memory bandwidth-bound. Model weights (10-1,000 GB) are fetched repeatedly from external memory during inference, as GPUs typically have only  100 MB of on-chip memory. For per request, generating thousands of tokens per second demands hundreds of TB/s bandwidth, far exceeding the capabilities of HBM (high bandwidth memory) on current GPUs.
While tensor parallelism across GPUs can increase bandwidth, mitigating communication bottlenecks in a large GPU cluster remains challenging. Also, adding GPUs improves throughput for concurrent queries but does not reduce time per output token (TPOT), as each query is still memory bandwidth-limited.
@@ 2.2 Reasons for wafer-scale accelerators
Area (mm2) Up to 858 Up to 73062
# Transistors (TSMC n3) ~1s Tillion ~10s Trillion
# Cores 1,000s-10,000s 100,000s-1,000,000s
On-Chip Memory 10s-100s MB ~10s GB
Memory Bandwidth 1s TB/s ~10s PBs/s
Attached HBM ~10s-100s GB 10s TB (via TSMC SoW) [21]
Die-to-Die
Bandwidth (TB/s)
~1s-10s (via off-chip) ~10s-100s (via on-chip)
Die-to-Die
Latency (ns)
~100s (via off-chip) ~1s (via on-chip)
Die-to-Die
Power (pJ/bit)
~10s (via off-chip) ~0.1s (via on-chip)
@@
Performance advantages. System-on-wafer technology allows trillions of transistors to be integrated into a single wafer-scale chip, up to 100$\times$ more than a typical GPU die, shown in Table 1. This enables millions of AI-optimized cores, providing tens of GBs of on-chip memory and up to tens of PB/s memory bandwidth, 1,000$\times$ higher than a standard GPU’s several TB/s. Future wafer-scale chips can also attach 40-80$\times$ more HBM chips to their edge compared to a standard die [21].
Integration efficiency. System-on-wafer excels at integrating massive parallel cores, with wafer-based die-to-die connections offering up to 10$\times$ more bandwidth per unit area, 100-300$\times$ latency benefit, and nearly 100$\times$ better power efficiency per bit than conventional PCB-based chip-to-chip interconnection [24] (e.g., NVIDIA NVLink, PCIe), shown in Table 1. As noted earlier, LLM inference is primarily constrained by memory bandwidth due to intensive data access. In distributed settings, inter-chip communication overhead from remote memory access, especially during decode-phase GEMV, limits scalability. Thus, system-on-wafer integration offers lower-latency and high-bandwidth interconnects and improved efficiency over conventional chip-to-chip links.
Lower cost. Wafer-scale integration can lower the manufacturing cost, since a significant fraction of the cost of fabrication (typically 30-50%) is related to testing and packaging the individual chips [48]. Additionally, wafer-scale integration has made notable progress in yield improvement. Companies such as TSMC are also developing techniques to integrate fully tested dies on a single wafer, further enhancing yield.
@@ 2.3 Challenges for wafer-scale LLM inference
The massive-scale mesh architecture presents challenges for several LLM operations due to their high data movement demands: (i) managing LLM models and KV cache, (ii) GEMM operations during the prefill phase, and (iii) GEMV operations during the decode phase. Other operations, such as element-wise computations such as dot-product and activation functions, require no data movement and naturally benefit from parallelism. Operations needing allreduce, such as RMSNorm and Softmax, can leverage GEMV solutions.
@@ 3.1 The PLMR model
    Massive Parallelism (P): A wafer-scale accelerator can accommodate millions of parallel cores, compared to thousands in GPUs. Each core features a local hardware pipeline that overlaps data ingress, egress, computation, and memory access at the cycle level. This requires the computation to be partitioned at a massive scale and a fine-grained schedule to overlap computation, memory access, and NoC communication.
@@
    Highly non-uniform memory access Latency (L): Accessing memory across cores in a mesh incurs highly non-uniform latency. In a mesh with $N_{w}\times N_{h}$ cores, organized as a rectangle, the maximum number of NoC hops between two cores is $N_{w}+N_{h}$, and the worst-case memory access latency is given by $\alpha(N_{w}+N_{h})+\beta r$, where $r<N_{w}+N_{h}$ denotes the number of routing stages along the communication path. Here, $\alpha$ denotes the per-hop transmission latency, the cost incurred when a message is directly forwarded at a core according to the router hardware’s pre-configured rules, increasing with hops. $\beta$ represents the per-routing latency, the overhead when a message is involved in header parsing and rewriting by software at a core when forwarding [27]. Typically, $\alpha<\beta$. In a mesh with a million cores, the maximum number of hops and routing can reach up to several thousand, resulting in up to a thousand times latency gap between local and remote memory access. Consequently, minimizing long-range communication is critical for performance.
@@
    Constrained per-core local Memory (M): Each core has a small local memory (tens of KBs to several MBs), as performance and energy efficiency decline with larger capacities [47]. As a result, computation data must be explicitly partitioned into fine-grained chunks to fully fit within the constraints of each core’s local memory.
@@
    Constrained Routing resources (R): Wafer-scale accelerators integrating millions of cores impose strict constraints on each core’s routing circuit complexity or routing table size. For example, on the Cerebras WSE-2, each core can recognize only message headers with a 5-bit address code. Consequently, each core can support at most $2^{5}$ distinct routing paths, and the software system must carefully plan these paths. For long-distance remote communication, a pair of cores can consume routing resources to establish a direct routing path, incurring only $\alpha$ latency. However, if the number of routing paths exceeds hardware limits, messages must be relayed through multiple intermediate cores, introducing additional $\beta$ latency.